What is the value today of $3,100 per year, at a discount rate of 9 percent, if the first payment is received 5 years from today and the last payment is received 15 years from today? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
1
Expert's answer
2020-05-12T10:43:53-0400

"PV=FV\/(1+i)^n"

i-discount rate

n-year

PV-present value

FV-future value

"PV=\frac{3100}{(1+0.09)^5}+\frac{3100}{1.09^6}+\frac{3100}{1.09^7}+\frac{3100}{1.09 ^8}+\frac{3100}{1.09^9}+\frac{3100}{1.09^{10}}+\frac{3100}{1.09^{11}}+\frac{3100}{1.09^{12}}+\frac{3100}{1.09^{13}}+\frac{3100}{1.09^{14}}+\frac{3100}{1.09^{15}}=14,945"